1.   Add the binary numbers 11102 and 1112

1
Expert's answer
2021-08-16T13:58:03-0400

"\begin{matrix}
\underline {\,\,\,\,\,\,\,\begin{matrix}
1&1&1&0\\
 + &{}&{}&{}\\
{}&1&1&1
\end{matrix}} \\
\begin{matrix}
1&0&1&0&1
\end{matrix}
\end{matrix}"

Answer: "{1110_2} + {111_2} = {10101_2}"